# Supplemetary Material for the paper

### v1

@@ Appendix C Supporting Lemmas and the Proof of the Main Theorem
holds until $t\in\left[0,T\right]$ or $\|(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t))\|\geq\tilde{D}$ where, $\tilde{D}=N^{\frac{1}{3}}\sqrt{2L+1}\left(\frac{D}{2}+0.5\right)^{\frac{1}{3}}$ and $D\geq 2\rho\sqrt{K}$.
Note that,
$(N(2L+1)+2K)\text{exp}\left(2\beta T\right)L\alpha_{V}^{2}$ $\overset{a}{\leq}(N(2L+1)+2K)\frac{1}{N\left({2L}+1\right)}$
$\leq 1+\frac{2K}{N\left({2L}+1\right)}\leq 1+\frac{2}{{2L}+1}\leq 2.$ (234)
The inequality (a) is obtained by substituting the upper bound on the value $\alpha_{V}$ from equation (204). The following inequalities follow by noting that $K\leq N$, and $L\geq 1$. Applying the reverse triangle inequality to (232) and using (234) yields $\|({\mathbf{g}}(t),{\mathbf{V}}(t))\|<\tilde{D}+2$.
Using Lemma C.3, we can translate the bound on the parameters to a bound on the matrices. This yields:
$\Big\|\mathbf{X}(t)\left.-\tilde{\mathbf{X}}(t)\right\|_{F}$ $<2(\tilde{D}+2)^{2}\sqrt{2(N(2L+1)+2K)}\text{exp}\left(\beta t\right)\alpha_{V}$ (235)
$<2(\tilde{D}+2)^{2}\sqrt{2NL(2L+3)}\text{exp}\left(\beta T\right)\alpha_{V},$ (236)
where $\beta=16NL^{3/2}\left((N+1)\mu C^{4}+MB_{Y}C\right)$ as derived in C.2. Let the degree of approximation be denoted as $\epsilon_{app}\in(0,1)$. To ensure this in (236) is less than a desired approximation error $\epsilon_{app}$, we introduce a stricter condition on the initialization as:
@@
$\alpha_{V}\leq\frac{\epsilon_{app}}{2(\tilde{D}+2)^{2}}\frac{\text{exp}(-\beta T)}{\sqrt{2LN\left({2L}+3\right)}}.$ (237)
Thus for $t\in\left[0,T\right]$ we have
$\left\|\mathbf{X}(t)-\tilde{\mathbf{X}}(t)\right\|_{F}$ $<\epsilon_{app}.$ (238)
The proof is complete if we can show that the condition $\left\|\mathbf{X}(t)-\tilde{\mathbf{X}}(t)\right\|_{F}<\epsilon_{app}$ holds for all $t\in[0,\text{min}\{T,T_{D}\}]$, where $T_{D}:=\inf\{t\geq 0\mid\|\mathbf{X}(t)\|_{F}\geq D\}$. Let $t^{\prime}\in\left[0,T\right]$ be the first instance at which
$\|(\tilde{\mathbf{g}}(t^{\prime}),\tilde{\mathbf{V}}(t^{\prime}))\|^{2}=\tilde{D}^{2}$ (239)
Using the definition of distance equation, we have:
$L\|(\tilde{\mathbf{g}}(t^{\prime})\|_{2}^{2}+\|\tilde{\mathbf{V}}(t^{\prime}))\|_{F}^{2}$ $=\tilde{D}^{2}.$ (240)
The proof is complete if we can show that for $t\in[0,t^{\prime})$ the bound (238) holds, at $t=t^{\prime}$, $\|\mathbf{X}(t^{\prime})\|\geq D$.
Assuming perfect balancedness condition ($\frac{1}{2}\|\mathbf{g}(t^{\prime})\|_{2}^{2}=\|\mathbf{V}(t^{\prime})\|_{F}^{2}$) in (240), we have:
$\|(\tilde{\mathbf{g}}(t^{\prime})\|_{2}^{2}$ $=2\tilde{D}^{2}/(2L+1).$ (242)
Recall that norm of the reference matrix $\tilde{\mathbf{X}}(t)$ is:
$\left\|\tilde{\mathbf{X}}(t^{\prime})\right\|_{F}^{2}$ $=\sum\limits_{n\in[N]}\tilde{g}_{n}^{4}(t^{\prime})\sum\limits_{l\in[L]}\tilde{V}_{ij}(t^{\prime})^{2}.$ (243)
From row-balancedness condition $\left(\frac{1}{2}\tilde{g}_{i}^{2}(t)=\sum\limits_{l\in[L]}\tilde{V}_{ij}^{2}(t)\right)$, we can write it as,
$\left\|\tilde{\mathbf{X}}(t^{\prime})\right\|_{F}^{2}$ $=\frac{1}{2}\sum\limits_{n\in[N]}\tilde{g}_{n}^{6}(t^{\prime}),$ (244)
Using the generalized mean inequality, namely, for $p<q$, $\left(\frac{1}{N}\sum\limits_{i=1}^{N}x^{p}\right)^{1/p}\leq\left(\frac{1}{N}\sum\limits_{i=1}^{N}x^{q}\right)^{1/q}$.
Let $p=1$ and $q=3$, $x=\tilde{g}_{n}^{2}(t^{\prime})$.
@@
$\left(\frac{1}{N}\sum\limits_{i}^{N}\tilde{g}_{n}^{2}(t^{\prime})\right)$ $\leq\left(\frac{1}{N}\sum\limits_{i}^{N}\left(\tilde{g}_{n}^{2}(t^{\prime})\right)^{3}\right)^{1/3}$ (246)
@@
$\left(\frac{1}{N}\sum\limits_{i}^{N}\tilde{g}_{n}^{2}(t^{\prime})\right)^{3}$ $\leq\left(\frac{1}{N}\sum\limits_{i}^{N}\left(\tilde{g}_{n}^{2}(t^{\prime})\right)^{3}\right)$ (247)
We can now substitute this into (244):
@@
$\left\|\tilde{\mathbf{X}}(t^{\prime})\right\|_{F}^{2}$ $\geq\frac{1}{2N^{2}}\left(\sum\limits_{n\in[N]}\tilde{g}_{n}^{2}(t^{\prime})\right)^{3}$ (248)
$\geq\frac{1}{2N^{2}}\left(\frac{2\tilde{D}^{2}}{2L+1}\right)^{3}$ (249)
$\geq\frac{4\tilde{D}^{6}}{((2L+1)^{3}N^{2})}.$ (250)
Substituting $\tilde{D}$ in (250) gives:
@@
$\left\|\tilde{\mathbf{X}}(t^{\prime})\right\|_{F}$ $\geq\frac{2\tilde{D}^{3}}{N(2L+1)^{3/2}}\geq D+1.$ (251)
But at time $t=0$, by the choice of our initialization we have $\|\tilde{\mathbf{X}}(0)\|_{F}\ =2\rho\sqrt{K}<D+1$. This implies that $t^{\prime}>0$. Now, we focus on the time interval $t\in[0,t^{\prime})$. Note that until $t<t^{\prime}$, $\left\|\mathbf{X}(t)-\tilde{\mathbf{X}}(t)\right\|_{F}\leq\epsilon_{app}<1$ holds.
Using continuity with respect to time, at $t=t^{\prime}$ we have
$\left\|\mathbf{X}(t^{\prime})-\tilde{\mathbf{X}}(t^{\prime})\right\|_{F}\leq\epsilon_{app}<1.$ (252)
Together with equation (251), and applying reverse triangular inequality,
$\left\|\mathbf{X}(t^{\prime})\right\|_{F}$ $\geq\left\|\tilde{\mathbf{X}}(t^{\prime})\right\|_{F}-\epsilon_{app}$ (253)
@@
$>\left\|\tilde{\mathbf{X}}(t^{\prime})\right\|_{F}-1$ (254)
$>D.$ (255)
This shows that at time $t=t^{\prime}$, $\left\|\mathbf{X}(t^{\prime})\right\|_{F}>D$. This guarantees that
$\left\|\mathbf{X}(t^{\prime})-\tilde{\mathbf{X}}(t^{\prime})\right\|_{F}\leq\epsilon_{app}$ (256)
until time $T$ or time $t^{\prime}$ at which $\left\|\mathbf{X}(t^{\prime})\right\|_{F}\geq D$. This completes the proof.
C.2 Proof of Corollary 1
Corollary C.0.1 (Corollary 5.1.1 restated).

### v2

@@ C.1.2 Bounding the error on support indices $(i\in\mathcal{S}_{K})$
Thus the analysis holds for $t\in[0,T)$ where, $T=\min\left\{\frac{2}{C_{\Lambda}^{\prime}\sqrt{\mathcal{E}_{e_{i}}(0)}},\frac{2}{\tilde{C}_{\Lambda}^{\prime}\sqrt{\mathcal{E}_{r_{i}}(0)}}\right\}$.
Provided $t\in[0,T)$, we can rewrite the lower bound (199) as
[…]
We now simplify the lower bound derived in (203). A closer inspection of the terms reveals a specific algebraic structure: the expression resembles the derivative of an inverse product, i.e., $\frac{\mathrm{d}}{\mathrm{d}t}\frac{-1}{h(t)f(t)}$. To exploit this, let us define two auxiliary functions $h(t)=\mathcal{E}_{e_{i}}^{-1/2}(0)-\frac{1}{2}C_{\Lambda}^{\prime}t$ and $f(t)=\mathcal{E}_{r_{i}}^{-1/2}(0)-\frac{1}{2}\tilde{C}_{\Lambda}^{\prime}t$.
Then $\frac{\mathrm{d}}{\mathrm{d}t}f(t)=-\frac{1}{2}\tilde{C}_{\Lambda}^{\prime}$ and $\frac{\mathrm{d}}{\mathrm{d}t}h(t)=-\frac{1}{2}C_{\Lambda}^{\prime}$. Recall that $\tilde{C}_{\Lambda}^{\prime}:=4\sqrt{2}\Bigg(L+\frac{1}{2}\Bigg)^{-\frac{1}{2}}\tilde{C}_{\Lambda}$ and $C_{\Lambda}^{\prime}:=4\sqrt{2}\Bigg(L+\frac{1}{2}\Bigg)^{-\frac{1}{2}}C_{\Lambda}$. Substituting these equations back into equation (203) and simplifying we get:
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}_{x_{i}}(t)$ $\geq\frac{h(t)\frac{\mathrm{d}}{\mathrm{d}t}f(t)+f(t)\frac{\mathrm{d}}{\mathrm{d}t}h(t)}{h^{2}(t)f^{2}(t)}$ (204)
Rewriting the term as the derivative of an inverse product:
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}_{x_{i}}(t)$ $\geq\frac{\mathrm{d}}{\mathrm{d}t}\frac{-1}{h(t)f(t)}$ (205)
Integrating both sides with respect to time from 0 to $t$
$\int\limits_{0}^{t}\frac{\mathrm{d}}{\mathrm{d}\tau}\mathcal{E}_{x_{i}}(\tau)\mathrm{d}\tau$ $\geq\int\limits_{0}^{t}\frac{\mathrm{d}}{\mathrm{d}\tau}\frac{-1}{h(\tau)f(\tau)}\mathrm{d}\tau$ (206)
$\mathcal{E}_{x_{i}}(t)-\mathcal{E}_{x_{i}}(0)$ $\geq\frac{1}{h(0)f(0)}-\frac{1}{h(t)f(t)}$ (207)
Recall that at $\mathcal{E}_{x_{i}}(0)=\sqrt{L}(2L+1)\alpha_{V}\rho^{1/3}$, $h(0)=\mathcal{E}_{e_{i}}^{-1/2}(0)=\left(\alpha_{V}\sqrt{L(2L+1)}\right)^{-1}$ and $f(0)=\mathcal{E}_{r_{i}}^{-1/2}(0)=\left(\sqrt{2L+1}\rho^{1/3}\right)^{-1}$.
@@
$\mathcal{E}_{x_{i}}(t)$ $\geq 2\sqrt{L}(2L+1)\alpha_{V}\rho^{1/3}-\frac{1}{\left(\mathcal{E}_{e_{i}}^{-1/2}(0)-\frac{1}{2}C_{\Lambda}^{\prime}t\right)\left(\mathcal{E}_{r_{i}}^{-1/2}(0)-\frac{1}{2}\tilde{C}_{\Lambda}^{\prime}t\right)}$ (208)
We now combine the bounds (173), (187) and (208) to obtain an upper bound on the error $\mathcal{E}_{i}(t)$ for all $i\in\mathcal{S}_{K}$. Recall from equation (174) that the total error decomposes as
$\mathcal{E}_{i}(t)$ $=\mathcal{E}_{e_{i}}(t)+\mathcal{E}_{r_{i}}(t)-2\mathcal{E}_{x_{i}}(t)$ (174)
Substituting the bounds derived from equations (173), (187) and (208):
@@
$\leq\frac{1}{\Big(\mathcal{E}_{e_{i}}^{-1/2}(0)-\frac{1}{2}C_{\Lambda}^{\prime}t\Big)^{2}}+\frac{1}{\Big(\mathcal{E}_{r_{i}}^{-1/2}(0)-\frac{1}{2}\tilde{C}_{\Lambda}^{\prime}t\Big)^{2}}+\frac{2}{\left(\mathcal{E}_{e_{i}}^{-1/2}(0)-\frac{1}{2}C_{\Lambda}^{\prime}t\right)\left(\mathcal{E}_{r_{i}}^{-1/2}(0)-\frac{1}{2}\tilde{C}_{\Lambda}^{\prime}t\right)}-4\sqrt{L}(2L+1)\alpha_{V}\rho^{1/3}$ (209)
Observe that the first three terms form perfect square:
@@
$\leq\left(\frac{1}{\mathcal{E}_{e_{i}}^{-1/2}(0)-\frac{1}{2}C_{\Lambda}^{\prime}t}+\frac{1}{\mathcal{E}_{r_{i}}^{-1/2}(0)-\frac{1}{2}\tilde{C}_{\Lambda}^{\prime}t}\right)^{2}-4\sqrt{L}(2L+1)\alpha_{V}\rho^{1/3}$ (210)
Since the initialization parameters $\alpha_{V}$ and $\rho$ are positive, we can omit this negative term to obtain a simplified upper bound:
@@
$\mathcal{E}_{i}(t)=\sum\limits_{l\in[L]}\Big(\big(g_{i}(t)-\tilde{g}_{i}(t)\big)^{2}+\big(V_{il}(t)-\tilde{V}_{il}(t)\big)^{2}\Big)$ $\leq\left(\frac{1}{\mathcal{E}_{e_{i}}^{-1/2}(0)-\frac{1}{2}C_{\Lambda}^{\prime}t}+\frac{1}{\mathcal{E}_{r_{i}}^{-1/2}(0)-\frac{1}{2}\tilde{C}_{\Lambda}^{\prime}t}\right)^{2}$ (211)
C.1.3 Bounding the total error in the parameter-space
Summing over all rows (both support and non-support) allows us to bound the total parameter distance squared, $\|(\mathbf{g}(t),\mathbf{V}(t))-(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t))\|^{2}$:
$\Big\|(\mathbf{g}(t),\mathbf{V}(t))-(\tilde{\mathbf{g}}(t),$ $\tilde{\mathbf{V}}(t))\Big\|^{2}$
$=\sum\limits_{i\in[N]}\sum\limits_{l\in[L]}\Big(\big|g_{i}(t)-\tilde{g}_{i}(t)\big|^{2}+\big|V_{il}(t)-\tilde{V}_{il}(t)\big|^{2}\Big)$
$=\sum\limits_{i\in[N]-\mathcal{S}_{K}}\sum\limits_{l\in[L]}\Big(\big|g_{i}(t)-\tilde{g}_{i}(t)\big|^{2}+\big|V_{il}(t)-\tilde{V}_{il}(t)\big|^{2}\Big)+\sum\limits_{i\in\mathcal{S}_{K}}\sum\limits_{l\in[L]}\Big(\big|g_{i}(t)-\tilde{g}_{i}(t)\big|^{2}+\big|V_{il}(t)-\tilde{V}_{il}(t)\big|^{2}\Big)$
$=(N-K)\mathcal{E}_{e_{i}}(t)+K\mathcal{E}_{i}(t)$ (212)
$\leq\frac{N-K}{\Big(\mathcal{E}_{e_{i}}^{-1/2}(0)-\frac{1}{2}C_{\Lambda}^{\prime}t\Big)^{2}}+K\left(\frac{1}{\mathcal{E}_{e_{i}}^{-1/2}(0)-\frac{1}{2}C_{\Lambda}^{\prime}t}+\frac{1}{\mathcal{E}_{r_{i}}^{-1/2}(0)-\frac{1}{2}\tilde{C}_{\Lambda}^{\prime}t}\right)^{2}$ (213)
Note that $\frac{1}{\left(\mathcal{E}_{e_{i}}^{-1/2}(0)-\frac{1}{2}C_{\Lambda}^{\prime}t\right)^{2}}<\left(\frac{1}{\mathcal{E}_{e_{i}}^{-1/2}(0)-\frac{1}{2}C_{\Lambda}^{\prime}t}+\frac{1}{\mathcal{E}_{r_{i}}^{-1/2}(0)-\frac{1}{2}\tilde{C}_{\Lambda}^{\prime}t}\right)^{2}$ since all terms are strictly positive. Using this inequality, we can upper bound the error contribution from the non-support rows (NKN-K terms) by the strictly larger squared sum expression used for the support rows. Furthermore, since NK<NN-K<N, we can combine the coefficients to simplify the total bound:
$\leq N\left(\frac{1}{\mathcal{E}_{e_{i}}^{-1/2}(0)-\frac{1}{2}C_{\Lambda}^{\prime}t}+\frac{1}{\mathcal{E}_{r_{i}}^{-1/2}(0)-\frac{1}{2}\tilde{C}_{\Lambda}^{\prime}t}\right)^{2}$ (214)
C.2 From Parameter Error to Signal Reconstruction Error
Finally, to guarantee the recovery of the signal matrix $\mathbf{X}$, we translate the error in parameter space into the Frobenius norm of the difference of the estimated trajectory and reference trajectory $\left(\Big\|\mathbf{X}(t)\left.-\tilde{\mathbf{X}}(t)\right\|_{F}^{2}\right)$. Utilizing the result from Lemma C.2, which relates parameter distance to matrix distance for bounded trajectories, we obtain:
$\Big\|\mathbf{X}(t)\left.-\tilde{\mathbf{X}}(t)\right\|_{F}^{2}$ $\leq 8D^{4}\Big\|(\mathbf{g}(t),\mathbf{V}(t))\left.-(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t))\right\|^{2}$ (215)
$\leq 8D^{4}N\left(\frac{1}{\mathcal{E}_{e_{i}}^{-1/2}(0)-\frac{1}{2}C_{\Lambda}^{\prime}t}+\frac{1}{\mathcal{E}_{r_{i}}^{-1/2}(0)-\frac{1}{2}\tilde{C}_{\Lambda}^{\prime}t}\right)^{2}$ (216)
@@ C.2 From Parameter Error to Signal Reconstruction Error
$\Big\|\mathbf{X}(t)\left.-\tilde{\mathbf{X}}(t)\right\|_{F}$ $\leq 2\sqrt{2N}D^{2}\left(\frac{1}{\mathcal{E}_{e_{i}}^{-1/2}(0)-\frac{1}{2}C_{\Lambda}^{\prime}t}+\frac{1}{\mathcal{E}_{r_{i}}^{-1/2}(0)-\frac{1}{2}\tilde{C}_{\Lambda}^{\prime}t}\right)$ (217)
Note that the right hand side of the equation (217) is monotonically increasing with time $t$. To ensure that the error $\Big\|\mathbf{X}(t)\left.-\tilde{\mathbf{X}}(t)\right\|_{F}$ remains bounded by $\epsilon_{\text{app}}$ until $t=T$, we restrict the analysis to a time interval $t\in[0,T)$ where the bound $\Big\|\mathbf{X}(t)\left.-\tilde{\mathbf{X}}(t)\right\|_{F}\leq\epsilon_{\text{app}}$ holds.
[…]
Let us define the worst case constants as $d(0):=\min\big\{\mathcal{E}_{e_{i}}(0)^{-1/2},\mathcal{E}_{r_{i}}(0)^{-1/2}\big\}$ and $C_{\max}=\max\big\{C_{\Lambda}^{\prime},\tilde{C}_{\Lambda}^{\prime}\big\}$. Then, at any $t<T$, where $T=\min\left\{\frac{2}{C_{\Lambda}^{\prime}\sqrt{\mathcal{E}_{e_{i}}(0)}},\frac{2}{\tilde{C}_{\Lambda}^{\prime}\sqrt{\mathcal{E}_{r_{i}}(0)}}\right\}$, the bound (218) simplifies to
$\frac{4\sqrt{2N}D^{2}}{d(0)-\frac{1}{2}C_{\max}T}\leq\epsilon_{\text{app}}$ (219)
Let $\rho$ be set such that the initial energy of estimated trajectory is at-least that of the reference trajectory:
$\mathcal{E}_{e_{i}}(0)$ $\geq\mathcal{E}_{r_{i}}(0)$ (220)
${L(2L+1)}\alpha_{V}^{2}$ $\geq(2L+1)\rho^{2/3}$ (221)
$L^{3/2}\alpha_{V}^{3}\geq\rho$ (222)
Under this setting, $d(0)=\mathcal{E}_{e_{i}}(0)^{-1/2}=\left({\alpha_{V}\sqrt{L(2L+1)}}\right)^{-1}$.
Note that
$C_{\max}$ $=\max\big\{C_{\Lambda}^{\prime},\tilde{C}_{\Lambda}^{\prime}\big\}$ (223)
@@
Substituting the constants $C_{\Lambda}^{\prime}=\frac{8}{\sqrt{2L+1}}\left(\|\mathbf{Y}\|_{F}+2L^{3/2}N\alpha_{V}^{3}\right),\quad\tilde{C}_{\Lambda}^{\prime}=\frac{8}{\sqrt{2L+1}}\left(\|\mathbf{Y}\|_{F}+\sqrt{K}N\rho\right)$
$C_{\max}$ $=\frac{8}{\sqrt{2L+1}}\max\bigg\{\|\mathbf{Y}\|_{F}+2L^{3/2}N\alpha_{V}^{3},\|\mathbf{Y}\|_{F}+\sqrt{K}N\rho\bigg\}$ (224)
$=\frac{8}{\sqrt{2L+1}}\cdot\left(\|\mathbf{Y}\|_{F}+N\max\bigg\{2L^{3/2}\alpha_{V}^{3},\sqrt{K}\rho\bigg\}\right)$ (225)
Substituting the bound (222)
$C_{\max}$ $\leq\frac{8}{\sqrt{2L+1}}\left(\|\mathbf{Y}\|_{F}+NL^{3/2}\alpha_{V}^{3}\max\bigg\{2,\sqrt{K}\bigg\}\right)$ (226)
To ensure that the approximation error $\epsilon_{\text{app}}$ remains small enough i.e., $\epsilon_{\text{app}}\in(0,1)$, time $T$ should satisfy equation (219).
Rewriting equation (219), and substituting the constants,
$\frac{4\sqrt{2N}D^{2}}{d(0)-\frac{1}{2}C_{\max}T}$ $<\epsilon_{\text{app}}$ (227)
Rearranging the terms:
$\frac{4\sqrt{2N}D^{2}}{\epsilon_{\text{app}}}$ $<d(0)-\frac{1}{2}C_{\max}T$ (228)
Grouping the constants on the right hand side and time dependent term on the other:
$\frac{1}{2}C_{\max}T$ $<d(0)-\frac{4\sqrt{2N}D^{2}}{\epsilon_{\text{app}}}$ (229)
Solving for $T$:
$T$ $<2\frac{d(0)-\frac{4\sqrt{2N}D^{2}}{\epsilon_{\text{app}}}}{C_{\max}}$ (230)
Simplifying the numerator and denominator:
$T$ $<2\frac{d(0)\epsilon_{\text{app}}-{4\sqrt{2N}D^{2}}}{\epsilon_{\text{app}}C_{\max}}$ (231)
Substituting constants $d(0)=\left({\alpha_{V}\sqrt{L(2L+1)}}\right)^{-1},C_{\max}=\frac{8}{\sqrt{2L+1}}\left(\|\mathbf{Y}\|_{F}+NL^{3/2}\alpha_{V}^{3}\max\bigg\{2,\sqrt{K}\bigg\}\right)$
$T$ $<2\frac{\frac{\epsilon_{\text{app}}}{\alpha_{V}\sqrt{L(2L+1)}}-{4\sqrt{2N}D^{2}}}{\frac{8}{\sqrt{2L+1}}\epsilon_{\text{app}}\left(\|\mathbf{Y}\|_{F}+NL^{3/2}\alpha_{V}^{3}\max\bigg\{2,\sqrt{K}\bigg\}\right)}$ (232)
Simplifying, we have
$T$ $<\frac{\epsilon_{\text{app}}-\alpha_{V}{4\sqrt{2NL(2L+1)}D^{2}}}{4\epsilon_{\text{app}}\sqrt{L}\alpha_{V}\left(\|\mathbf{Y}\|_{F}+NL^{3/2}\alpha_{V}^{3}\max\bigg\{2,\sqrt{K}\bigg\}\right)}$ (233)
To ensure that the upper bound on $T$ is positive, the numerator should be positive. This limits $\alpha_{V}$ as follows:
$0$ $<\epsilon_{\text{app}}-\alpha_{V}4\sqrt{2NL(2L+1)}D^{2}$ (234)
$\alpha_{V}4\sqrt{2NL(2L+1)}D^{2}$ $<\epsilon_{\text{app}}$ (235)
$\alpha_{V}$ $<\frac{\epsilon_{\text{app}}}{4\sqrt{2NL(2L+1)}D^{2}}$ (236)
This initialization ensures that the error $\|\mathbf{X}(t)-\tilde{\mathbf{X}}(t)\|_{F}$ remains below $\epsilon_{app}$ for any time $t\in[0,T)$.
C.3 Proof of Corollary 1
Corollary C.3 (Corollary 5.7 restated).